The hypotenuse of a right triangle is 6 more than the shorter leg. The longer leg is three more than the shorter leg. Find the length of the shorter leg.
Answer by mananth(16946)

The hypotenuse of a right triangle is 6 more than the shorter leg. The longer leg is three more than the shorter leg. Find the length of the shorter leg.
Let the shorter leg be x
hypotenuse = x+6
Longer leg is x+3
(x+6)^2 = x^2+ (x+3)^2
x^2+12x+36 = x^2+x^2+6x+9
x^2-6x-27 =0
x^2-9x+3x-27=0
x(x-9)+3(x-9)
(x+3)(x-9)=0
x=9
the sides are 9, 12 15